Help please
Problem 24.28 A thin, horizontal, 19-cm-diameter copper plate is charged to -2.6 nC. Assume that the electrons are uniformly distributed on the surface
media%2F186%2F186e9856-677f-4a08-b45a-12
media%2F059%2F059dd7ff-eac7-4b26-9291-a0
media%2F4cb%2F4cbadc1e-e495-4fa2-873b-c0
Answer #1

Part A

x=0.1 mm= 0.1*10^(-3) m

R=9.5 cm=0.095 m

=Q/pi*R^2= -9.2*10^(-8) C/m^2

we got

E=5197 N/C

Part C

at center x=0

E=5202.5 N/C

Part D

E=5197 N/C